OpenStudy (amking):

Which expression is equivalent to cot t sec t? csc t sin t tan t sin2 t csc2 t

OpenStudy (mathstudent55):

\(\Large \cot t \sec t\) \(\Large = \dfrac{\cos t}{\sin t} \times \dfrac{1}{\cos t} \) Can you simplify the rest?